a paper company makes rectangular pieces of cardboard for packing computers. the sheet of cardboard are to have an area of 54 square inches., and the length of the sheet is to be 5 inches more that 2/3 the width. determine the length and width of the cardboard to be manufactured.
Answer by Cromlix(4381)

Width = x
Length = (5 + 2/3x)
Area = Length * Width
54 = (5 + 2/3x) * x
54 = 5x + 2/3x^2
2/3x^2 + 5x - 54 = 0
(2/3x + 9)(x - 6) = 0
x = -27/2 (No answer)
OR
x = 6
Width = 6 inches
Length = 9 inches (5 + 2/3*6)
6 * 9 = 54 inches^2